# COOCO - Common Objects Out-of-Context - Semantic Violation in Scenes: Investigating Multimodal Context in Referential Communication

### v1

@@ 9 Conclusion
Repeated exposure to common visual scenes generates expectations about what objects are likely to occur (the ‘semantics’ of the scene; Võ, 2021). In this paper, we address whether state-of-the-art multimodal models acquire such knowledge in the course of training. Experiments on COOCO, a novel dataset which carefully controls for the target-scene congruence, reveals that scene semantics play an important role in the way in which models generate references to objects, though the extent to which models rely on scene context versus the target region itself is modulated by perturbations to these image regions. Our findings are consistent with recent work on simple, artificial scenes, which suggest that VLMs leverage scene context Junker and Zarrieß (2024, 2025). Our analysis of attention is also consistent with findings from a growing body of interpretability research, which suggests that middle layers of VLMs play a crucial role in cross-modal fusion Zhang et al. (2025a, b); Golovanevsky et al. (2025).
@@ Acknowledgements
This research was supported by the Netherlands Science Foundation (NWO) under the AiNed AiNed XS Europe 2023, grant number NGF.1609.23.020. FM was a visiting student from the University of Trento to Utrecht University during the period when the work was conducted. We thank Sina Zarrieß, Simeon Junker and Hendrik Buschmeier for their valuable comments and helpful discussions.
Appendix A Further details on COOCO image creation
living room, arrival gate outdoor, restaurant, market/outdoor, parking lot, bakery shop, arrival gate/outdoor, art studio, gas station, banquet hall, tower, street, music studio, fastfood restaurant, pantry, cubicle/office, market outdoor, kitchen, coffee shop, bedroom, dorm room, cubicle office, delicatessen, train railway, bakery/shop, home office, physics laboratory, dining room, bathroom
@@ Appendix A Further details on COOCO image creation
A.1 Description of the inpainting and verification pipeline
For inpainting, we use the LaMa inpainting model Suvorov et al. (2021) to remove the target object from a scene, expanding its mask by 20% to make sure that the resulting image is free of irregularities. We then create a square-padded version of the cleaned image, and create a mask for object inpainting,
For each of the three relatedness groups (low, medium, and high), we retain 15 replacement candidates. In the high-similarity condition only, the similarity to the original target object is also considered. Additionally, we generate one image containing the same target object as in the original image to use as a control condition. See Figure 9 for the distributon of relatedness scores per relatedness level.
@@ A.1 Description of the inpainting and verification pipeline
Appendix B Spatial Encoding & Model Prompts
@@
xGen-MM xGen-MM embeds bounding box coordinates within the prompt using <bbox> tags and learns to interpret these as visual region references in natural language form. Prompt: <|user|> <image>What is the object in this part <bbox>[x1, y1][x2, y2]</bbox> of the image? Answer with the object’s name only. No extra text.<|end|> <|assistant|>
@@
Qwen2.5-VL builds upon the Qwen-VL architecture by integrating visual tokens into a large language model using enhanced spatial encoding strategies. It employs bounding box references enclosed in square bracket format $[x_{1},y_{1},x_{2},y_{2}]$, where coordinates are normalized within the range [0,1]. The spatial references are directly inserted into the natural language prompt, making them interpretable as explicit multimodal instructions.
Qwen2.5-VL Qwen2.5-VL encodes normalized bounding box coordinates directly in the prompt as $[x_{1},y_{1},x_{2},y_{2}]$, supporting explicit multimodal alignment through instruction-tuned sequences. Prompt: What is the object in this part of the image [x1, y1, x2, y2]? Answer with the object’s name only. No extra text.
@@
Appendix C Attention over image tokens in LLaVA-OneVision
@@
    All visual tokens are projected via a two-layer MLP into the LLM’s embedding space;
@@
    The resulting image tokens are prepended to the tokenized question together with an answer-start marker;
@@
In Section 7, we restrict our analysis to the visual tokens obtained from the fixed encoding of the full image, omitting any multi-crop tokens.
Below, we detail the process of aggregating attention values and computing the layerwise attention allocation ratio $r^{(t)}_{l}$ described in Section 7.
C.1 Aggregating LLM Attention
@@
C.2 Aggregating Attention in Vision Transformers (ViTs)
During our initial experiments, we noticed a tendency where the model applied attention over a specific set of tokens for many images, indicating a model-intrinsic bias. To mitigate such systematic attention biases, we compute a layer-wise average attention map across the dataset and subtract it from each instance’s attention map. This adjustment emphasizes instance-specific patterns by removing model-intrinsic tendencies. To avoid resolution-based confounds, we include only images of size 640$\times$640 (which comprise 86% of the dataset).
@@ C.2 Aggregating Attention in Vision Transformers (ViTs)
C.3 Attention Over Image Regions
@@
C.4 Attention Over Target and Context
@@
C.5 Attention Allocation Ratio
To analyze attention allocation across different levels $l$ of the vision encoder and experimental instances $t$, we compute the ratio $r$ between the total normalized attention values assigned to the target ($a_{\mathrm{target}}$) and the context ($a_{\mathrm{context}}$): $r_{l}^{(t)}=\frac{a_{\mathrm{target},l}^{(t)}}{a_{\mathrm{context},l}^{(t)}}$. A value of $r$ closer to 1 indicates a higher allocation of attention to the target region, whereas a value closer to 0 suggests greater attention to the context. Since the context area is larger than the target, we do not expect $r$ to exceed 0.5. Instead, our focus is on analyzing its variations across different experimental conditions.

### v2

@@ 8 Conclusion
This paper presented COOCo, a dataset to explore scene-semantic violations. Consistent with recent work Junker and Zarrieß (2024); Junker and ZarrieSS (2025), we show that VLMs rely on scene semantics to refer to objects, but this is modulated by perturbations to these image regions. While these results hold across different models, and are supported by in-depth analysis of LLaVA-OneVision-0.5B, we believe that detailed comparison of different architectures would shed further light on the role of context in VLMs. We replicated our findings on VISIONS Allegretti et al. (2025), which differs from COOCo in that it does not involve artificial image manipulation.
Our analysis of attention under different conditions contributes to a growing body of interpretability research aimed to localise cross-modal fusion in VLMs Zhang et al. (2025a, b); Golovanevsky et al. (2025).
@@ Acknowledgements
This research was supported by the Netherlands Science Foundation (NWO) under the AiNed XS Europe 2023 program, grant number NGF.1609.23.020. FM was a visiting student from the University of Trento to Utrecht University during the period when the work was conducted. We thank Sina ZarrieSS, Simeon Junker and Hendrik Buschmeier for their valuable comments and helpful discussions. We also thank our four anonymous TACL reviewers and the action editor, whose input helped to improve the paper.
Appendix A Additional preprocessing, data and model details
A.1 COOCo image creation
living room, arrival gate outdoor, restaurant, market/outdoor, parking lot, bakery shop, arrival gate/outdoor, art studio, gas station, banquet hall, tower, street, music studio, fastfood restaurant, pantry, cubicle/office, market outdoor, kitchen, coffee shop, bedroom, dorm room, cubicle office, delicatessen, train railway, bakery/shop, home office, physics laboratory, dining room, bathroom.
@@ A.1 COOCo image creation
A.1.1 Inpainting and verification pipeline
We use the LaMa inpainting model Suvorov et al. (2021) to remove the target object from a scene, expanding its mask by 20% to make sure that the resulting image is free of irregularities. We then create a square-padded version of the cleaned image, and create a mask for object inpainting, For each of the three relatedness groups (low, medium, and high), we retain 15 replacement candidates. In the high-similarity condition only, the similarity to the original target object is also considered. Additionally, we generate one image containing the same target object as in the original image to use as a control condition. See Figure 9 for the distributon of relatedness scores per relatedness level.
@@ A.1.1 Inpainting and verification pipeline
A.2 Spatial encoding & model prompts
@@
xGen-$M$ xGen-$M$ embeds bounding box coordinates within the prompt using <bbox> tags and learns to interpret these as visual region references in natural language form. Prompt: <|user|> <image>What is the object in this part <bbox>[x1, y1][x2, y2]</bbox> of the image? Answer with the object’s name only. No extra text.<|end|> <|assistant|>
@@
Qwen2.5-VL builds upon the Qwen-VL architecture by integrating visual tokens into a large language model using enhanced spatial encoding strategies. It employs bounding box references enclosed in square bracket format $x_{1},y_{1},x_{2},y_{2}]$, where coordinates are normalized within the range [0,1]. The spatial references are directly inserted into the natural language prompt, making them interpretable as explicit multimodal instructions.
Qwen2.5-VL Qwen2.5-VL encodes normalized bounding box coordinates directly in the prompt as $x_{1},y_{1},x_{2},y_{2}]$, supporting explicit multimodal alignment through instruction-tuned sequences. Prompt: What is the object in this part of the image [x1, y1, x2, y2]? Answer with the object’s name only. No extra text.
@@
A.3 Accuracy comparison between dataset labels and modal object names in VISIONS
Table 4: Contingency table showing the counts and row percentages of agreement and disagreement between accuracy computed with dataset labels (rows) and with modal names (columns), along with the overall correlation and number of divergent cases.
Counts
Accuracy 0 1
             0 4817 73
Accuracy Modal Names
             1 95 955
Row Percentages
Accuracy 0 1
             0 98.1 7.1
Accuracy Modal Names
             1 1.9 92.9
Pearson correlation: 0.902
Divergent cases: 168 / 5940 (2.83%)
We use the modal names provided in the VISIONS dataset to compare our cosine-similarity-based accuracy metric under two different ground truths: (1) the dataset labels and (2) the human-generated modal names. This comparison allows us to assess how sensitive the metric is to variation in naming. The contingency tables 4 show a strong alignment between the two ground-truth conditions. Predictions judged incorrect using the dataset label are almost always incorrect using the modal name as well (98.1%), and predictions judged correct with the dataset label are likewise correct with the modal name in most cases (92.9%). The low divergence rate (2.83%) and high correlation (0.902) indicate that disagreements between the two ground-truth settings are rare and that the metric captures nearly identical performance patterns across them. Overall, these results suggest that the accuracy metric is robust to synonym variation and is not substantially affected by differences in naming between dataset labels and human-provided modal names.
A.4 Attention over image tokens in LLaVA-OneVision
@@
    All visual tokens are projected via a two‐layer MLP into the LLM’s embedding space;
@@
    The resulting image tokens are prepended to the tokenized question together with an answer‐start marker;
@@
In Section 6.2, we restrict our analysis to the visual tokens obtained from the fixed encoding of the full image, omitting any multi‐crop tokens.
Below, we detail the process of aggregating attention values and computing the layerwise attention allocation ratio $r^{(t)}_{l}$ described in Section 6.2.
A.5 Aggregating LLM Attention
@@
A.6 Aggregating Attention in Vision Transformers (ViTs)
During our initial experiments, we noticed a tendency where the model applied attention over a specific set of tokens for many images, indicating a model-intrinsic bias. To mitigate such systematic attention biases, we compute a layer-wise average attention map across the dataset and subtract it from each instance’s attention map. This adjustment emphasizes instance-specific patterns by removing model-intrinsic tendencies. To avoid resolution-based confounds, we include only images of size $\times640$ (which comprise 86% of the dataset).
@@ A.6 Aggregating Attention in Vision Transformers (ViTs)
A.7 Attention Over Image Regions
@@
A.8 Attention Over Target and Context
@@
A.9 Attention Allocation Ratio
To analyze attention allocation across different levels $l$ of the vision encoder and experimental instances $t$, we compute the ratio $r$ between the total normalized attention values assigned to the target ($a_{\mathrm{target}}$) and the context ($a_{\mathrm{context}}$): $r_{l}^{(t)}=\frac{a_{\mathrm{target},l}^{(t)}}{a_{\mathrm{context},l}^{(t)}}$. A value of $r$ closer to 1 indicates a higher allocation of attention to the target region; a value closer to 0 suggests greater attention to the context. Since the context area is larger than the target, we do not expect $r$ to exceed 0.5. Instead, our focus is on analyzing its variations across different experimental conditions.
@@ A.9 Attention Allocation Ratio
Appendix B Complementary results
Congruent images Incongruent images
Coef. Std.Err. z Coef. Std.Err. z
Int. 1.744 0.054 32.406 1.745 0.055 31.902
Context -0.087 0.013 -6.56 -0.178 0.013 -13.518
Target -0.105 0.013 -7.866 -0.094 0.013 -7.152
N. level -0.040 0.009 -4.261 -0.028 0.009 -2.953
N. level $\times$ Con 0.009 0.013 0.667 -0.001 0.013 0.044
N. level $\times$ Tar 0.041 0.013 3.030 0.023 0.013 1.78
Group Var 0.020 0.044 0.020 0.046
Table 5: Linear mixed effects model on VISIONS: Noise area (Context/Target) and noise level (N. level) on scene-target semantic similarity, for congruent and incongruent conditions. denotes a significant effect at $p\leq.01$. Conditional $R^{2}$ values: 0.423 (congruent) and 0.452 (incongruent).
Figure 10: Mean semantic similarity between model outputs and scene labels as a function of binned continuous object–scene relatedness (COOCo) and object consistency (VISIONS), using 10 bins. Columns correspond to datasets; rows to noise–area conditions (target, context, all). Continuous lines represent the baseline (0.0 noise) for correct (blue) and incorrect responses (orange); dotted lines indicate the 0.5-noise condition; shaded regions show the variability introduced when increasing noise from 0.5 to 1.0.
(a) COOCo, $\Delta\bar{r}_{l}$ (range: $\pm 0.25$)
(b) VISIONS, $\Delta\bar{r}_{l}$ (range: $\pm 0.32$).
Figure 11: Attention-to-target delta $\Delta\bar{r}_{l}$, between correct and incorrect predictions across layers, semantic relatedness levels, noise levels, and noise areas. Heatmaps are predominantly green, indicating that when the model correctly recognizes the object, it tends to allocate more attention to the target token compared to when it misclassifies the object.
COOCo dataset
0.5 noise – Target
0.5 noise – Context
0.5 noise – All
1 noise – Target
1 noise – Context
1 noise – All
VISIONS dataset
0.5 noise – Target
0.5 noise – Context
0.5 noise – All
1 noise – Target
1 noise – Context
1 noise – All
Figure 12: Mean attention-to-target ratio as a function of semantic fit for COOCo and VISIONS, at two noise levels (0.5, 1.0) and noise area conditions (target, context, all).
Table 6: Accuracy results for LLaVA-OneVision-0.5B on the COOCo dataset across relatedness levels, noise intensities, and noise areas.
Rel. Level Noise Level Noise Area Total Samples Correct Samples Accuracy (%)
low 0.0 111384 41860 37.58
0.5 all 111384 10816 9.71
context 111384 41548 37.30
target 111384 4732 4.25
1.0 all 111384 3302 2.96
context 111384 41990 37.70
target 111384 728 0.65
medium 0.0 117494 44954 38.26
0.5 all 117494 17810 15.16
context 117494 45448 38.68
target 117494 11674 9.94
1.0 all 117494 6084 5.18
context 117494 43134 36.71
target 117494 3510 2.99
high 0.0 34086 17654 51.79
0.5 all 34086 8242 24.18
context 34086 17758 52.10
target 34086 5252 15.41
1.0 all 34086 2938 8.62
context 34086 16692 48.97
target 34086 1820 5.34
same target 0.0 40326 34476 85.49
0.5 all 40326 23452 58.16
context 40326 34736 86.14
target 40326 21008 52.10
1.0 all 40326 10192 25.27
context 40326 33722 83.62
target 40326 8970 22.24
original 0.0 41626 34840 83.70
0.5 all 41626 19734 47.41
context 41626 37024 88.94
target 41626 18434 44.28
1.0 all 41626 8320 19.99
context 41626 35256 84.70
target 41626 9490 22.80
Table 7: Accuracy results for LLaVA-OneVision-0.5B on the VISIONS dataset across congruency conditions, noise intensities, and noise areas for incongruent and congruent conditions.
Rel. Level Noise Level Noise Area Total Samples Correct Samples Accuracy (%)
Incongruent 0.0 8580 1950 22.73
0.5 All 8580 104 1.21
Context 8580 1742 20.30
Target 8580 52 0.61
1.0 All 8580 104 1.21
Context 8580 1898 22.12
Target 8580 52 0.61
Congruent 0.0 8580 3250 37.88
0.5 All 8580 390 4.55
Context 8580 2626 30.61
Target 8580 624 7.27
1.0 All 8580 78 0.91
Context 8580 2834 33.03
Target 8580 624 7.27